OpenStudy (anonymous):

Use the definition of the derivative to find the derivative of f(x)=6

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

0

OpenStudy (saifoo.khan):

Totally zero

OpenStudy (anonymous):

rate of a constant =0

OpenStudy (anonymous):

How do I know its zero? how do you find that?

OpenStudy (anonymous):

becuase the derivative of any constant is 0 ...like 6 or 7 or 5000000

OpenStudy (anonymous):

what if F(x)=2x^3-1

OpenStudy (anonymous):

6x^2

OpenStudy (anonymous):

how do u know that?

OpenStudy (anonymous):

you need to calculate: \[\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}\] knowing that:\[f(x)=6\] so \[f(x+h)=6\]

OpenStudy (anonymous):

that is the definition of a derivative.

OpenStudy (anonymous):

brought the exponent down and multiplied it to the coefficient then subtracted the exponent by one

OpenStudy (anonymous):

it seems like the poster is just learning the rules of the derivative, so posting rules that we all know about derivatives isnt what he/she is looking for.

OpenStudy (anonymous):

yes ive just started learning derivatives

OpenStudy (anonymous):

Since f(x) = 6, is the constant function, f(anything) = 6. So when calculating the limit we obtain: \[\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}=\lim_{h\rightarrow 0}\frac{6-6}{h}=\lim_{h\rightarrow 0}\frac{0}{h}=0\]

OpenStudy (anonymous):

You need to do it from first principals lim(f(x+h) -f(x))/h as h --> 0 lim(6 - 6)/h = 0/h = 0 fx = 6 f(x+h) = 6

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

if it was f(x) = x +6 it would be lim((x +h +6) - x +6)/h lim(h/h) = 1